# Faces of homogeneous cones and applications to homogeneous chordality

### v2

@@ 3. $(iii)$
where the first equality follows from (2.1). The second equality holds because $Q_{u}$ is an automorphism, so it maps a face to another face and is also a bijection between the relative interiors. Finally, the last equality holds because $e_{I}\in\mathrm{ri}\,(\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I}))={\mathcal{K}}(\mathcal{A}_{I})$. We note that if $|I|=0$ we have $e_{I}=e$ and ${\mathcal{K}}(\mathcal{A}_{I})={\mathcal{K}(\mathcal{A})}$, which would contradict the fact that ${\mathcal{F}}$ is proper. We have thus obtained item $(ii)$ since homogeneity and rank are preserved by linear isomorphisms (see Proposition 2.4) and ${\mathcal{K}}(\mathcal{A}_{I})$ is a homogeneous cone of rank $r-|I|<r$.
@@
=${\mathcal{F}}=\{x\in\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}\mid\langle Q_{u}(x),\bar{e}\rangle=0\}=\{x\in\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}\mid\langle x,Q_{u^{*}}(\bar{e})\rangle=0\},$ (3.7)
@@
Let ${\mathcal{F}}=\mathrm{Face}(x,\mathrm{cl}\,{\mathcal{K}(\mathcal{A})})$. We observe that $x$ generates an extreme ray if and only if the dimension of ${\mathcal{F}}$ is one. By Theorem 3.6, ${\mathcal{F}}$ is linearly isomorphic to $\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$, where $I=\{i\mid\rho_{i}(t_{ii})=0\}$. Since $e_{i}\in\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$ for $i\not\in I$, the dimension of $\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$ is at least $r-|I|$. So the only way that ${\mathcal{F}}$ can be an extreme ray is if $r-|I|=1$, which happens if and only if $\rho_{i}(t_{ii})>0$ for one and exactly one $i\in\{1,\ldots,r\}$. Conversely, if $r-|I|=1$ then a direct verification shows that the dimension of $\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$ is one, so ${\mathcal{F}}$ is an extreme ray. ∎
@@
Recalling that $e_{I}=\sum_{i\not\in I}e_{i}$, we have that $(ue_{I})_{ij}=\sum_{k=1}^{r}u_{ik}(e_{I})_{kj}$ is 00 if $j\in I$ and $u_{ij}$ if $j\not\in I$. Since $ue_{I}=e_{J}$ must hold, we have $u_{ij}=0$ for $i\neq j$, i.e., $u$ only has nonzero components on its “diagonal”. Finally, since $u\in\mathcal{T}_{++}$, $ue_{I}=e_{J}$ forces that $I=J$. ∎
@@ 3.3 The dual cone
A potential pitfall is the definition of proper (lower) triangular matrices. Applying the definition of proper (upper) triangular matrices to $\mathcal{B}$ and translating the indices back to $\mathcal{A}$, we see the “correct way” of defining properness. Indeed, $t\in\mathcal{L}_{+}(\mathcal{A})$ is said to be proper if $\rho_{i}(t_{ii})=0$ implies that the $i$-th “column” (not the row!) of $t$ vanishes [^3] This is somewhat confusing, so here is a quick example. For $r=3$, let $t$ be an upper triangular element in $\mathcal{B}$ and let us look at the third column, which corresponds to $t_{13},t_{23},t_{33}$. In view of (2.7), if we decompose $t$ in the bigradation of $\mathcal{A}$ (i.e., express $t$ as sum of elements in the $\mathcal{A}_{ij}$’s), $t$ corresponds to a lower triangular $\hat{t}\in\mathcal{L}(\mathcal{A})$ and the third column of $t$ gets mapped to the first column of $\hat{t}$, so that $t_{33}=\hat{t}_{11}$, $t_{23}=\hat{t}_{21}$ and $t_{13}=\hat{t}_{31}$. . Analogous to Algorithm 1, we also have a Cholesky factorization in proper lower triangular matrices. Essentially, we apply Algorithm 1 to $\mathcal{B}$ and translate back the indices to $\mathcal{A}$ in two steps. First, the indices in $t_{ij}$ and $y_{ij}$ are replaced with $r+1-i$ and $r+1-j$, respectively and, then we simplify the resulting for loop. The result is Algorithm 2.
@@ 2. $(ii)$
    ${\mathcal{F}}$ is a homogeneous cone of rank $s<r$ and there is a triangular automorphism $Q_{l}$ of ${\mathcal{K}(\mathcal{A})}^{*}$ such that $Q_{l}({\mathcal{F}})=\mathrm{cl}\,({\mathcal{K}}^{D}(\mathcal{A}_{I}))$ holds for some $I\subsetneq\{1,\ldots,r\}$. Furthermore, letting $x\in\mathrm{ri}\,{\mathcal{F}}$ and writing $x=tt^{*}$ for a proper $t\in\mathcal{L}_{+}(\mathcal{A})$, we have $I=\{i\mid\rho_{i}(t_{ii})=0\}$ and $s=r-|I|$.
@@ 3. $(iii)$
Let ${\mathcal{K}(\mathcal{A})}$ be a homogeneous cone of rank $r$ and let ${\mathcal{F}}\mathrel{\unlhd}\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$ be a face of rank $s$. Then, ${\mathcal{F}}^{\Delta}={\mathcal{K}(\mathcal{A})}^{*}\cap{\mathcal{F}}^{\perp}$ has rank $r-s$.
@@
We have that $e_{I}=\sum_{i\not\in I}e_{i}$ and $|I|=r-s$. $Q_{u^{*}}$ is a triangular automorphism of ${\mathcal{K}(\mathcal{A})}^{*}$, so $Q_{u^{*}}(y)\in{\mathcal{K}(\mathcal{A})}^{*}$ if and only if $y\in{\mathcal{K}(\mathcal{A})}^{*}$. In addition, as the $e_{i}$’s are in $\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$, for $y\in{\mathcal{K}(\mathcal{A})}^{*}$ we have
@@
That is, ${\mathcal{K}(\mathcal{A})}^{*}\cap\{x\}^{\perp}$ is linearly isomorphic to a principal dual face of rank $r-|J|=|I|=r-s$, which concludes the proof. ∎

### v3

@@ 3. $(iii)$
where the first equality follows from (2.1). The second equality holds because $Q_{u}$ is an automorphism, so it maps a face to another face and is also a bijection between the relative interiors. Finally, the last equality holds because $e_{I}\in\mathrm{ri}\,(\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I}))={\mathcal{K}}(\mathcal{A}_{I})$. We note that if $|I|=0$ we have $e_{I}=e$ and ${\mathcal{K}}(\mathcal{A}_{I})={\mathcal{K}(\mathcal{A})}$, which would contradict the fact that $\mathcal{F}$} is proper. We have thus obtained item $(ii)$ since homogeneity and rank are preserved by linear isomorphisms (see Proposition 2.4) and ${\mathcal{K}}(\mathcal{A}_{I})$ is a homogeneous cone of rank r|I|<rr-|I|<r.
@@
${\mathcal{F}}=\{x\in\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}\mid\langle Q_{u}(x),\bar{e}\rangle=0\}=\{x\in\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}\mid\langle x,Q_{u^{*}}(\bar{e})\rangle=0\},$ (3.7)
@@
Finally, we remark that Chua had previously observed a description of the faces of the PSD cone using triangular matrices, see [Chu06, Equation (2.4)]. Chua also proved the following result for homogeneous cones. Let $y\in{\mathcal{F}}$ and define $B$ to be the set of $i$’s for which there exists at least one $z\in{\mathcal{F}}$ whose unique Cholesky factor $t$ satisfies $t_{ii}\neq 0$. Writing $y=\hat{t}\hat{t}^{*}$ for a proper matrix $\hat{t}$, Chua showed that $y\in\mathrm{ri}\,{\mathcal{F}}$ if and only if $\rho_{i}(\hat{t})>0$ for all $i\in B$, see [Chu06, Proposition 18]. Intuitively this means that, inside $\mathcal{F}$}, the elements of the relative interior are precisely the ones that have as many nonzeros as possible in the diagonal of their unique Cholesky factors. We note that $B$ itself is not enough decide membership on $\mathrm{ri}\,{\mathcal{F}}$ for an arbitrary $y\in\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$, since there are finitely many possible $B$’s and $\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$ may have infinitely many faces. Nevertheless, under the setting of Theorem 3.6, we have $I=\{1,\ldots,r\}\setminus B$, so $B$ identifies the principal face of $\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$ that is linearly isomorphic to $\mathcal{F}$}. Although the statement [Chu06, Proposition 18] is only about membership in the relative interior of $\mathcal{F}$} for some $y\in{\mathcal{F}}$, its actual proof hints at an expression for $\mathcal{F}$} that is somewhat related to item $(ii)$ of Theorem 3.6.
Before we move on, we present an example.
Example 3.7.
We consider the matrix norm cone, which is a spectrahedral cone used in rank minimization heuristics, see [TV23, Section 5.1]. It was analyzed in slightly different form in [GT98, Example 5]. We define it as
$\mathcal{M}_{m,n}\coloneqq\left\{x\in\mathcal{S}_{+}^{m+n}\mid x=\begin{pmatrix}V&U\\ U^{*}&\alpha I_{n}\end{pmatrix},\alpha\in\mathbb{R},U\in\mathbb{R}^{m\times n},V\in\mathbb{R}^{m\times m}\right\},$
where $I_{n}$ is the $n\times n$ identity matrix and $U^{*}$ is the usual matrix tranpose of $U$.
We consider the following algebra of matrices
$\mathcal{A}^{m,n}\coloneqq\left\{x\in\mathbb{R}^{(m+n)\times(m+n)}\mid x=\begin{pmatrix}V&W\\ U^{*}&\alpha I_{n}\end{pmatrix},\alpha\in\mathbb{R},W\in\mathbb{R}^{m\times n},U\in\mathbb{R}^{m\times n},V\in\mathbb{R}^{m\times m}\right\}.$ (3.8)
The multiplication is given by
$\begin{pmatrix}V_{1}&W_{1}\\ U_{1}^{*}&\alpha_{1}I_{n}\end{pmatrix}\begin{pmatrix}V_{2}&W_{2}\\ U_{2}^{*}&\alpha_{2}I_{n}\end{pmatrix}=\begin{pmatrix}V_{1}V_{2}+W_{1}U_{2}^{*}&V_{1}W_{2}+\alpha_{2}W_{1}\\ U_{1}^{*}V_{2}+\alpha_{1}U_{2}^{*}&(\textup{tr}(U_{1}W_{2}^{*})+\alpha_{1}\alpha_{2})I_{n}\end{pmatrix}$,
where $\textup{tr}(U_{1}W_{2}^{*})$ is the usual matrix trace of $U_{1}W_{2}^{*}$. The multiplication in $\mathcal{A}^{m,n}$ differs from the usual matrix multiplication only in the lower right block.
It turns out that $\mathcal{A}^{m,n}$ is a $T$-algebra of rank $m+1$. Let $x$ be as in (3.8) and we will describe the bigradation of $\mathcal{A}^{m,n}$ in terms of the components of $x$. $\mathcal{A}^{m,n}_{m+1,m+1}$ is isomorphic to $\mathbb{R}$ and corresponds to $\alpha$, i.e., $\mathcal{A}^{m,n}_{m+1,m+1}\coloneqq\left\{\begin{pmatrix}0&0\\ 0&\alpha I_{n}\end{pmatrix}\mid\alpha\in\mathbb{R}\right\}$. For $i<m+1$, $\mathcal{A}^{m,n}_{i,m+1}$ is isomorphic to $\mathbb{R}^{n}$ and it corresponds to the $i$-th row of $W$, while $\mathcal{A}^{m,n}_{m+1,i}$ is isomorphic to $\mathbb{R}^{n}$ and it corresponds to the $i$-th column of $U^{*}$. Finally, for i,j<m1i,j<m-1, $\mathcal{A}^{m,n}_{i,j}$ corresponds to the $(i,j)$-entry of $V$. The involution is the usual matrix transpose. We have $e_{m+1}=\begin{pmatrix}0&0\\ 0&I_{n}\end{pmatrix}$ and, for $i<m+1$, $e_{i}$ is the diagonal matrix having $1$ in the $(i,i)$-entry and zero elsewhere. The T-algebra trace function $\textup{tr}(x)$ satisfies $\textup{tr}(x)=\textup{tr}(V)+\alpha$, where $\textup{tr}(V)$ is the usual matrix trace of $V$.
With that, axioms 1 and 2 of T-algebra definition are satisfied. Checking axioms 3-7 is cumbersome but relatively straightforward given that the multiplication only differs from the usual matrix multiplication in computations involving the block $\mathcal{A}^{m,n}_{m+1,m+1}$. For this case, it is more convenient to check 6 and 7 by verifying, equivalently, that for upper triangular $t,u,w\in\mathcal{T}(\mathcal{A}^{m,n})$ we have $t(uw)=(tu)w$ and $t(uu^{*})=(tu)u^{*}$ respectively, see [Chu09, Remarks 6 and 7]. In particular, 6 is immediate given that the multiplication in $\mathcal{A}^{m,n}$ coincides with the usual matrix multiplication for upper triangular matrices.
This formulation allow us to apply Theorem 3.6 to $\mathcal{M}_{m,n}$. Every principal subalgebra $\mathcal{A}^{m,k}_{I}$ is either a usual matrix algebra, if $m+1\in I$, or a smaller $\mathcal{A}^{k,n}$ with $k\leq m$ if $m+1\not\in I$. This implies that all faces of $\mathcal{M}_{m,n}$ are either linearly isomorphic to $\mathcal{S}_{+}^{k}$ or to $\mathcal{M}_{k,n}$ for $k\leq m$. Furthermore, $\mathcal{M}_{m,n}$ is projectionally exposed.
@@
Let ${\mathcal{F}}=\mathrm{Face}(x,\mathrm{cl}\,{\mathcal{K}(\mathcal{A})})$. We observe that $x$ generates an extreme ray if and only if the dimension of $\mathcal{F}$} is one. By Theorem 3.6, $\mathcal{F}$} is linearly isomorphic to $\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$, where $I=\{i\mid\rho_{i}(t_{ii})=0\}$. Since $e_{i}\in\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$ for $i\not\in I$, the dimension of $\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$ is at least r|I|r-|I|. So the only way that $\mathcal{F}$} can be an extreme ray is if r|I|=1r-|I|=1, which happens if and only if $\rho_{i}(t_{ii})>0$ for one and exactly one $i\in\{1,\ldots,r\}$. Conversely, if r|I|=1r-|I|=1 then a direct verification shows that the dimension of $\mathrm{cl}\,{\mathcal{K}}(\mathcal{A}_{I})$ is one, so $\mathcal{F}$} is an extreme ray. ∎
@@
Recalling that $e_{I}=\sum_{i\not\in I}e_{i}$, we have that $(ue_{I})_{ij}=\sum_{k=1}^{r}u_{ik}(e_{I})_{kj}$ is 0 if $j\in I$ and $u_{ij}$ if $j\not\in I$. Since $ue_{I}=e_{J}$ must hold, we have $u_{ij}=0$ for $i\neq j$, i.e., $u$ only has nonzero components on its “diagonal”. Finally, since $u\in\mathcal{T}_{++}$, $ue_{I}=e_{J}$ forces that $I=J$. ∎
@@ 3.3 The dual cone
A potential pitfall is the definition of proper (lower) triangular matrices. Applying the definition of proper (upper) triangular matrices to $\mathcal{B}$ and translating the indices back to $\mathcal{A}$, we see the “correct way” of defining properness. Indeed, $t\in\mathcal{L}_{+}(\mathcal{A})$ is said to be proper if $\rho_{i}(t_{ii})=0$ implies that the $i$-th “column” (not the row!) of $t$ vanishes [^3] This is somewhat confusing, so here is a quick example. For $r=3$, let $t$ be an upper triangular element in $\mathcal{B}$ and let us look at the third column, which corresponds to $t_{13},t_{23},t_{33}$. In view of (2.7), if we decompose $t$ in the bigradation of $\mathcal{A}$ (i.e., express $t$ as sum of elements in the $\mathcal{A}_{ij}$’s), $t$ corresponds to a lower triangular $\hat{t}\in\mathcal{L}(\mathcal{A})$ and the third column of $t$ gets mapped to the first column of $\hat{t}$, so that $t_{33}=\hat{t}_{11}$, $t_{23}=\hat{t}_{21}$ and $t_{13}=\hat{t}_{31}$. . Analogous to Algorithm 1, we also have a Cholesky factorization in proper lower triangular matrices. Essentially, we apply Algorithm 1 to $\mathcal{B}$ and translate back the indices to $\mathcal{A}$ in two steps. First, the indices in $t_{ij}$ and $y_{ij}$ are replaced with r+1ir+1-i and r+1jr+1-j, respectively and, then we simplify the resulting for loop. The result is Algorithm 2.
@@ 2. $(ii)$
    $\mathcal{F}$} is a homogeneous cone of rank $s<r$ and there is a triangular automorphism $Q_{l}$ of $\mathcal{K}(\mathcal{A})$$^{*}$ such that $Q_{l}({\mathcal{F}})=\mathrm{cl}\,({\mathcal{K}}^{D}(\mathcal{A}_{I}))$ holds for some nonempty $I\subseteq\{1,\ldots,r\}$. Furthermore, letting $x\in\mathrm{ri}\,{\mathcal{F}}$ and writing $x=tt^{*}$ for a proper $t\in\mathcal{L}_{+}(\mathcal{A})$, we have $I=\{i\mid\rho_{i}(t_{ii})=0\}$ and s=r|I|s=r-|I|.
@@ 3. $(iii)$
Let $\mathcal{K}(\mathcal{A})$} be a homogeneous cone of rank $r$ and let ${\mathcal{F}}\mathrel{\unlhd}\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$ be a face of rank $s$. Then, ${\mathcal{F}}^{\Delta}={\mathcal{K}(\mathcal{A})}^{*}\cap{\mathcal{F}}^{\perp}$ has rank rsr-s.
@@
We have that $e_{I}=\sum_{i\not\in I}e_{i}$ and |I|=rs|I|=r-s. $Q_{u^{*}}$ is a triangular automorphism of $\mathcal{K}(\mathcal{A})$$^{*}$, so $Q_{u^{*}}(y)\in{\mathcal{K}(\mathcal{A})}^{*}$ if and only if $y\in{\mathcal{K}(\mathcal{A})}^{*}$. In addition, as the $e_{i}$’s are in $\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$, for $y\in{\mathcal{K}(\mathcal{A})}^{*}$ we have
@@
That is, ${\mathcal{K}(\mathcal{A})}^{*}\cap\{x\}^{\perp}$ is linearly isomorphic to a principal dual face of rank r|J|=|I|=rsr-|J|=|I|=r-s, which concludes the proof. ∎
@@
3.4 Other approaches and related results
The T-algebra framework is not the only the tool suitable for the analysis of homogeneous cones. In this subsection, we discuss some other possibilities, related results and justify our usage of T-algebras.
Siegel cones
Let $t\in\mathbb{R},u\in\mathbb{R}^{n-1},x\in\mathcal{S}^{n-1}$. The matrix $\begin{pmatrix}t&u\\ u^{*}&x\end{pmatrix}$ is positive definite if and only if $t>0$ and its Schur complement $x-u^{*}u/t$ is positive definite. Or, equivalently, $t>0$ and $xt-u^{*}u$ is positive definite. Notably, $xt-u^{*}u$ being positive definite” is a constraint over $(n-1)\times(n-1)$ matrices. This allow us to construct $\mathcal{S}_{+}^{n}$ recursively via smaller dimensional PSD cones and we can repeat this argument until we reach $\mathcal{S}_{+}^{1}$, which is isomorphic to the nonnegative orthant. We can also view this construction in a bottom-up fashion and, starting from the nonnegative orthant, the cone $\mathcal{S}_{+}^{n}$ arises by $n$ applications of Schur complements.
A similar technique is available for homogeneous cones. Given a regular homogeneous cone ${\mathcal{K}}\subset\mathbb{R}^{n}$ and a bilinear form $B:\mathbb{R}^{p}\times\mathbb{R}^{p}\to\mathbb{R}^{n}$, the Siegel cone of $\mathcal{K}$} and $B$ is defined as
$\operatorname{SC}({\mathcal{K}},B)\coloneqq\mathrm{cl}\,\{(x,u,t)\in\mathbb{R}^{n}\times\mathbb{R}^{p}\times\mathbb{R}\mid t>0,~tx-B(u,u)\in{\mathcal{K}}\}$
and it is homogeneous if $B$ is a so-called “homogeneous $\mathcal{K}$}-bilinear form”, whose precise definition we omit. Importantly, every closed regular homogeneous cone $\mathcal{K}$} of dimension at least two is linearly isomorphic to a Siegel cone $\operatorname{SC}({\mathcal{K}}_{1},B_{1})$, where ${\mathcal{K}}_{1}$ is a lower dimensional homogeneous cone ${\mathcal{K}}_{1}$ and $B_{1}$ is a homogeneous ${\mathcal{K}}_{1}$-bilinear form (see [Rot66], [Gin92, Chapter 2], [GT98, Section 3]). In this way, starting from the nonnegative orthant, every closed regular homogeneous cone can be obtained by repeated applications of this construction.
T-algebras and Siegel cones are complementary approaches that generalize distinct aspects of PSD matrices. While T-algebras provide a generalized notion of Cholesky factorization, the Siegel cone construction generalizes Schur complements.
Naturally, arguments involving Siegel cones tend to be inductive. If we wish to show that all homogeneous cones have some property up to linear isomorphism, all we need to show is that the nonnegative orthant satisfies it (the easy part) and that if a homogeneous cone $\mathcal{K}$} satisfies it then $\operatorname{SC}({\mathcal{K}},B)$ satisfies it too (the delicate part). For instance, this is how the facial exposedness of homogeneous cones was established in [TT03]. Similarly, although we did not explore this possibility, it is likely possible to prove the projectional exposedness of homogeneous cones via an inductive argument using Siegel cones.
The Siegel cone approach allows very elegant proofs, but in our opinion, it is less convenient for describing a given arbitrary homogeneous cone. After all, while we may use Schur complements to prove properties of PSD matrices, we rarely think of a $n\times n$ positive definite matrix as $n$ Schur complements taken in sequence.
Similarly, for an arbitrary homogeneous cone $\mathcal{K}$}, it is useful to furnish the ambient space with an appropriate basis and reason about membership in $\mathcal{K}$} in terms of this basis in a direct (i.e., non-recursive) way, which is what the T-algebra framework gives us. Of course, it is not impossible to do something analogous under the Siegel cone framework. Doing so, however, naturally leads to algebras of generalized matrices that play a very similar role to the triangular matrices in the T-algebra construction, see, for example, [Gin92, pg.75-77 and §1.Theorem 2]. In our opinion, going along this route is not significantly different from using T-algebras from the beginning and has a similar amount of algebraic complexity.
We emphasize that rather than furnishing the shortest possible proofs of our statements, we are interested in convenient descriptions of the objects appearing in our discussion.
Matrix realizations
Let $\mathcal{K}$} be a homogeneous cone. Then $\mathcal{K}$} is spectrahedral, which means that it is linearly isomorphic to an intersection of the form $\mathcal{S}_{+}^{n}\cap\mathcal{L}$, where $\mathcal{L}$ is a subspace of the space of $n\times n$ real symmetric matrices $\mathcal{S}^{n}$. Here, we refer to $\mathcal{S}_{+}^{n}\cap\mathcal{L}$ as a matrix realization of $\mathcal{K}$}. As discussed in [TV23, Section 6], the existence of matrix realizations of homogeneous cones is implicit in earlier works by Vinberg [Vin63] and Rothaus [Rot66] which provided the theoretical backing for later approaches [Chu03, Fay02, Ish06, Ish15].
The geometric properties of an arbitrary spectrahedral cone can be quite different from homogeneous cones. For example, as mentioned previously, it is not known whether arbitrary spectrahedral cones are projectionally exposed. Therefore, as far as we can see, results such as Theorem 3.6 do not seem to follow simply from the mere existence of matrix realizations.
Nevertheless, discussions on matrix realizations of homogeneous cones often include results on how to obtain a group of linear transformations that act transitively into the relative interior of $\mathcal{S}_{+}^{n}\cap\mathcal{L}$, e.g., see [Ish15, Theorem 3]. With that, although we have not worked out the details, it seems entirely plausible to give alternative proofs of the results we have shown so far by making use of matrix realizations and appropriate descriptions of transitive subgroups of the automorphism group. One advantage of this approach is that it does not use T-algebras, thus bypassing some of the algebraic unpleasantness and bringing the discussion into the more familiar territory of PSD matrices.
A disadvantage is that matrix realizations may be wasteful because for an arbitrary homogeneous cone $\mathcal{K}$} of dimension $n$, in the worst case, we may need $n\times n$ PSD matrices to represent it, e.g., see [Chu03, Corollary 4.3] and [Ish11, Theorem 4]. This worst case is indeed achieved, e.g., for ${\mathcal{K}}=\mathbb{R}^{n}_{+}$ as we may verify that a necessary condition for $\mathcal{S}_{+}^{m}\cap\mathcal{L}$ to be linearly isomorphic to $\mathbb{R}^{n}_{+}$ is that $m\geq n$ holds [^4] The longest chain of faces of $\mathcal{S}_{+}^{m}$ is $m+1$ and the faces of $\mathcal{S}_{+}^{m}\cap\mathcal{L}$ are of the form $\mathcal{F}$$\cap\mathcal{L}$ where $\mathcal{F}$} is a face of $\mathcal{S}_{+}^{m}$ satisfying $(\mathrm{ri}\,{\mathcal{F}})\cap\mathcal{L}\neq\emptyset$. Furthermore, suppose that ${\mathcal{F}}_{1}$ and ${\mathcal{F}}_{2}$ are two such faces of $\mathcal{S}_{+}^{m}$ and that the proper inclusion ${\mathcal{F}}_{1}\cap\mathcal{L}\subsetneq{\mathcal{F}}_{2}\cap\mathcal{L}$ holds. Then, we have $\mathrm{ri}\,({\mathcal{F}}_{1}\cap\mathcal{L})=\mathrm{ri}\,({\mathcal{F}}_{1})\cap\mathcal{L}\subseteq{\mathcal{F}}_{2}\cap\mathcal{L}$, where the first equality follows from [Roc97, Corollary 6.5.1]. That is, $\mathrm{ri}\,({\mathcal{F}}_{1})$ passes through ${\mathcal{F}}_{2}$, which implies that ${\mathcal{F}}_{1}\subsetneq{\mathcal{F}}_{2}$. As the longest chain of faces of $\mathbb{R}^{n}_{+}$ is $n+1$, the only way we can fit a chain of faces of length $n+1$ into $\mathcal{S}_{+}^{m}\cap\mathcal{L}$ is if $m\geq n$ holds. .
When considering different formulations for the same optimization problem, if all else is the same, it is preferable to use the one for which the ambient space has dimension as small as possible. That is one of the reasons why it is rare to solve a second-order cone program by resorting to semidefinite programming formulations, see also related comments in [Chu03, Section 1].
Another undesirable fact is that a cone and its dual may need different $n$’s in their matrix realizations. For example the Vinberg cone and its dual are both homogeneous cones of rank $3$. While the latter has a natural matrix representation over $3\times 3$ matrices, the former cannot be written in the form $\mathcal{S}_{+}^{3}\cap\mathcal{L}$, see this footnote [^5] We only provide a sketch of the argument. The Vinberg cone is defined as ${\mathcal{K}}\coloneqq\left\{(y_{1},y_{2},y_{3},y_{4},y_{5})\in\mathbb{R}^{5}\mid\begin{pmatrix}y_{1}&y_{2}\\ y_{2}&y_{4}\end{pmatrix}\in\mathcal{S}_{+}^{2},\begin{pmatrix}y_{1}&y_{3}\\ y_{3}&y_{5}\end{pmatrix}\in\mathcal{S}_{+}^{2}\right\}$. Nonzero polynomials of minimal degree that vanish on the boundary of $\mathcal{K}$} must have degree $4$, see [GIL24, Section 4.1]. If $\mathcal{K}$} were linearly isomorphic to $\mathcal{S}_{+}^{3}\cap\mathcal{L}$ for some subspace $\mathcal{L}$, the restriction of the determinant function to $\mathcal{S}_{+}^{3}\cap\mathcal{L}$ would lead to a polynomial of degree at most $3$ that vanishes on the boundary of $\mathcal{K}$}, which is a contradiction. On the other hand, $\mathcal{K}$$^{*}$ is $\left\{(x_{1},x_{2},x_{3},x_{4},x_{5})\in\mathbb{R}^{5}\mid\begin{pmatrix}x_{1}&x_{2}&x_{3}\\ x_{2}&x_{4}&0\\ x_{3}&0&x_{5}\\ \end{pmatrix}\in\mathcal{S}_{+}^{3}\right\}$, which indeed has a matrix realization over $3\times 3$ matrices. Finally, we remark that whether $\mathcal{K}$} or $\mathcal{K}$$^{*}$ is referred as the “Vinberg cone” depends on the reference. Ishi calls $\mathcal{K}$} the “Vinberg cone” [Ish00, Section 5.2], while, up to a permutation of rows and columns, Tunçel and Vanderberghe call $\mathcal{K}$$^{*}$ the “Vinberg cone” [TV23, Example 6.3]..
The T-algebra representation, albeit cumbersome, is economical since both $\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$ and its dual $(\mathrm{cl}\,{\mathcal{K}(\mathcal{A})})^{*}$ are realized as cones of generalized matrices in the same space $\mathcal{H}(\mathcal{A})$, which has the same dimension of $\mathrm{cl}\,{\mathcal{K}(\mathcal{A})}$ and allows, for example, an intuitive discussion of the complementarity between ranks of conjugate faces as in Proposition 3.12.
Connection to Ishi’s results